1.) The mailroom is on the 6th floor. The mailman goes up 2 floors to deliver the mail. Next he goes up 7 floors. Then he goes down 11 floors. What floor is he on?
Thank yall so much

Answer by checkley77(12844)

6+2+7-11=15-11=4 the floor he is on (even if there is no 13 th. floor).